OpenStudy (anonymous):

solve by factoring j^2-17j+72+0

OpenStudy (dumbcow):

find 2 factors of 72 that add up to -17 they need to both be neg

OpenStudy (anonymous):

ok doin that now one sec

OpenStudy (anonymous):

i got neg 8 and 9

OpenStudy (dumbcow):

correct -8 * -9 = 72 -8 + -9 = -17 so j ^2-17j +72 = (j-8)(j-9) = 0 solve each binomial when equal to 0 j-8 = 0 and j-9 = 0

OpenStudy (anonymous):

hey thanks i think i got it from here but wil u check it when im done

OpenStudy (dumbcow):

sure